# Time-inversion of spatiotemporal beam dynamics using uncertainty-aware latent evolution reversal

### v1

@@ 2.3 Reverse latent evolution model
The forward discretized spatiotemporal beam dynamics problem can be written as $X_{m}=H(X_{1},X_{2},...,X_{m-1})$, where $H$ is a unknown nonlinear function that can be learned as $P(X_{M}|X_{1},X_{2},...,X_{M-1})$ with a latent evolution model [18]. Reversing the temporal dynamics to predict the current state $X_{m}$ given future states $X_{m+1},X_{m+2},\dots,X_{M}$, requires inverting the function $H$, which becomes a challenging inverse problem. Additionally, the time-varying nature of RF parameters and initial distribution further complicates the problem. The inverse problem can be learned by calculating the joint probability distribution $P(X_{1},X_{2},...,X_{M})$. This distribution can be computed with a reverse autoregression, where $P(X_{1},X_{2},...,X_{M})$ is factored as $P(X_{M})P(X_{M-1}|X_{M})P(X_{M-2}|X_{M-1},X_{M})$.
$P(X_{m}|X_{m+1},X_{m+2},\dots,X_{M})$.
For particle accelerators, $X$ is a high-dimensional object comprising positions and momentum of billions of particles. Learning such a high-dimensional probability distribution and calculating $P(X_{m}|X_{m+1},X_{m+2},\dots,X_{M})$ to estimate $X_{m}$ at any given module is computationally challenging. Fortunately, latent variable models (like variational autoencoders) transform the higher-dimensional distribution $P(X_{1},X_{2},...,X_{M})$ into a lower-dimensional distribution $P(\mathbf{z}_{1},\mathbf{z}_{2},...,\mathbf{z}_{M})$ using Bayesian inference [60]. The reverse autoregressive approach can then be applied to this lower-dimensional distribution, calculating $P(z_{m}|z_{m+1},z_{m+2},\dots,z_{M})$ to estimate $\mathbf{z}_{m}$. A recurrent neural network (RNN) like an LSTM can be used to invert the forward temporal dynamics in the latent space. The LSTM needs to be trained autoregressively to learn $\mathbf{z}_{m}=G(\mathbf{z}_{m-1:M})$ for all $m$, where $G$ is the combined functional relationships of LSTM [61]. This can be accomplished by training the LSTM with a variable input size. Once the inverse temporal dynamics is learned in the latent space, the decoder part of the VAE transforms it back to the high-dimensional space, obtaining $P(X_{m}|X_{m+1},X_{m+2},\dots,X_{M})$.
@@
The autoregressive procedure used to train the LSTM is a crucial aspect of rLEM. It is a general formulation which does not assume a strict Markov property, i.e. $P(z_{m}|z_{m+1},z_{m+2},\dots,z_{M})=P(z_{m}|z_{m+1})$, on the temporal dynamics. Since the temporal dynamics are learned in a lower-dimensional latent space rather than the higher-dimensional image space ($n_{d}\approx 10^{6}$), $M$-step autoregression compensates for the information loss due to spatial dimensionality reduction. A single step autoregression model, using $P(z_{m}|z_{m+1})$, suffers in performance compared to the $M$-step autoregressive model. Additionally, autoregression procedure leverages temporal periodicity (short-term or long-term seasonality) present in the dataset better than Markov property.
The reverse latent evolution approach described above combines a conditional variational autoencoder (CVAE) to project the 15 unique projections of the 6D phase space of charged particle beams into a lower-dimensional latent distribution. Subsequently, a LSTM network learns the reverse temporal dynamics within this latent space. During the prediction stage, both networks are integrated into an autoregressive loop to estimate all the upstream phase space projections given limited downstream measurements. The architecture of the rLEM is represented in Fig. 2. The model learns spatial and temporal correlations independently in two-step self-supervised process. The encoder performs convolution operations to extract features from the images. The module number is passed through dense layers and concatenated with the features vectors out of the encoder. The concatenated vector is the input to the latent space. The LSTM network operates on the learned latent space to predict the previous latent space points (corresponding to projections in the upstream modules) based on future points (corresponding to projections in the downstream modules). The predicted latent points are passed through a decoder which generates phase space projections across different modules of the accelerator.
@@
A mathematical description of the working of the rLEM algorithm can be written in terms of the CVAE and LSTM, combined through latent space $\mathbf{z}$. The posterior distribution $p_{\theta}(\mathbf{z}|\mathbf{x},m)$ (conditioned on the the module number in our work) can be written using Bayes rule:
$p_{\theta}(\mathbf{z}|\mathbf{x},m)=\frac{p_{\theta}(\mathbf{x}|\mathbf{z},m)p_{\theta}(\mathbf{z}|m)p_{\theta}(m)}{p_{\theta}(\mathbf{x})}.$ (4)
The above posterior is intractable and cannot be solved. CVAE introduces a parameterized distribution $q_{\phi}(\mathbf{z}|\mathbf{x},m)$ to reduce the dimensionality of the problem. It is an inference model that maps data $\mathbf{x}$ to the latent space $\mathbf{z}$ by approximating the posterior distribution $p_{\theta}(\mathbf{z}|\mathbf{x},m)$. Since, $p_{\theta}(m)$ is known and provides $p_{\theta}(\mathbf{z}|m)$, therefore, $p_{\theta}(\mathbf{x}|\mathbf{z},m)=p_{\theta}(\mathbf{x}|\mathbf{z})$. The loss function can be written in the form of Evidence Lower BOund (ELBO) [62]:
$ELBO(\theta,\phi;\mathbf{x})$ $=$ $\mathbb{E}_{\mathbf{z}\sim q_{\phi}(\mathbf{z}|\mathbf{x},m)}[\log p_{\theta}(\mathbf{x}|\mathbf{z})]$
$-D_{KL}(q_{\phi}(\mathbf{z}|\mathbf{x},m)||p_{\theta}(\mathbf{z}|m)).$
In ELBO, the first term represents the reconstruction loss between the original and the predicted data whereas the second term is the Kullback–Leibler (KL) divergence between the posterior distribution and the approximate posterior [63].
An LSTM is employed as a backcasting model to predict previous states (upstream modules) of the latent space based on limited downstream states. The LSTM equations [64], originally expressed in terms of input $x_{t}$ and output $y_{t}$ at time $t$, are reformulated to backcast $\mathbf{z}_{m-1}$ (output) given $\mathbf{z}_{m}$ (input), with $m$ denoting the module number:
$i_{m}=\sigma(W_{ii}\mathbf{z}_{m}+b_{ii}+W_{hi}h_{m+1}+b_{hi})$ (5)
$f_{m}=\sigma(W_{if}\mathbf{z}_{m}+b_{if}+W_{hf}h_{m+1}+b_{hf})$ (6)
$g_{m}=\tanh(W_{ig}\mathbf{z}_{m}+b_{ig}+W_{hg}h_{m+1}+b_{hg})$ (7)
$o_{m}=\sigma(W_{io}\mathbf{z}_{m}+b_{io}+W_{ho}h_{m+1}+b_{ho})$ (8)
$c_{m}=f_{m}\odot c_{m+1}+i_{m}\odot g_{m}$ (9)
$h_{m}=o_{m}\odot\tanh(c_{m})$ (10)
$\mathbf{z}_{m-1}=\sigma(W_{hy}h_{m}+b_{hy})$ (11)
where, $h_{m}$, $c_{m}$, $\mathbf{z}_{m}$ are the hidden state, cell state and input at module $m$. $i_{m}$, $f_{m}$, $g_{m}$, $o_{m}$ are respectively the input, forget gate, cell and output gates. $h_{m+1}$ is the hidden state at module $m+1$. $\sigma$ is the logistic sigmoid activation and $\odot$ is the Hadamard product. $W$ and $b$ are respective weights and bias matrices. A mean squared error-based loss function is used between true $\mathbf{z}_{m}$ and predicted $\hat{\mathbf{z}}_{m}$ values of latent point at module $m$.
$MSE=L(\psi;\mathbf{z})=\parallel\mathbf{z}_{m}-\hat{\mathbf{z}}_{m}\parallel_{2}$ (12)
The aforementioned equations are simplified with single input and single output but can be written multiple inputs and a single output.
@@
During the prediction phase, the trained CVAE and LSTM are integrated in an autoregressive loop to facilitate the predictions of phase space projections in upstream modules. The methodology is outlined in Algorithm 1. The input of the algorithm is the user-defined phase space at downstream location of the measurement ($m_{end}$) and the upstream module until where phase space predictions are desired, denoted by $m_{start}$. Typically, $m_{start}$ is set to 1, corresponding to the index of the first module, but it can also be any other module number $<m_{end}$. $m_{end}$ can be a single module, or multiple modules together. For the first iteration, the encoder of CVAE projects $X_{M:m}$ into the latent space to obtain $\mathbf{z}_{M:m}$. Subsequent iterations use only the LSTM and decoder of the CVAE. The LSTM acts on the latent representation to predict $\mathbf{z}_{m-1}$. The decoder of the CVAE then reconstructs $X_{m-1}$, providing the phase space projections at the previous module. The predicted latent point $\mathbf{z}_{m-1}$ is then concatenated with input $\mathbf{z}_{M:m}$ to form $\mathbf{z}_{M:m-1}$. This autoregressive procedure is iteratively applied to compute $\mathbf{z}_{m-2}$, followed by $X_{m-2}$, and so forth, until $m_{start}$.
Algorithm 1 rLEM prediction algorithm
1:Input: $m_{start}$, $m_{end}$, trained CVAE, trained LSTM
2:for $m$ = $m_{end}$rightarrow $m_{start}$ do
@@
4:         $\mathbf{z}_{M:m}$ = CVAE-encoder($X_{M:m}$, $M:m$)
5:     $\mathbf{z}_{m-1}$ = LSTM($\mathbf{z}_{M:m}$)
6:     $X_{m-1}$ = CVAE-decoder($\mathbf{z}_{m-1}$)
7:     $\mathbf{z}_{M:m-1}$ = concat($\mathbf{z}_{M:m}$, $\mathbf{z}_{m-1}$)
3 Results and Discussion
We have collected 1400 training data samples and 500 test data samples by running HPSim for different random RF settings. Every sample contains 15 unique phase space projections are depicted as 15 channels, each represented by a 256 $\times$ 256 pixel image along the 48 modules of the accelerator. The entire data set has size 1400 × 48 × 15 × 256 × 256 . The conditional input m to the encoder is the module number, a scalar between 1-48, normalized to the range [0, 1].
The dataset comprising 1400 data objects is partitioned into training and validation sets with an 85:15 ratio. The encoder of the CVAE is composed of five convolutional layers with 32, 64, 128, 256, and 512 filters of size 3×3, each with a stride of 2, followed by a dense layer with 256 neurons. Additionally, a separate dense branch processes the module number through two layers of 32 neurons. The outputs of the primary and branch layers are concatenated and fed into the latent dense layer. We have performed a trial-and-error study balancing dimensions and loss to select the dimensionality of the latent space. This lower-dimensional representation enhances interpretability and reduces training time. The decoder mirrors the encoder’s architecture, featuring a dense layer followed by convolutional layers with 512, 256, 128, 64, and 32 filters of size 3×3. Each layer employs the LeakyReLU activation function and batch normalization.
@@ 3 Results and Discussion
3.1 Latent space visualizations
@@ 3.2 Prediction results
The LSTM training with variable-length input and a single output allows the network to learn the autoregressive temporal patterns of the latent points, which is essential for making autoregressive predictions in the latent space. Our LSTM architecture comprises two layers, each with 64 hidden units. We minimize the mean squared error (MSE) between the true and predicted latent vectors, as defined in Eq. 12, using an Adam optimizer. The initial learning rate is set to 0.001 and is halved after 10 epochs if the validation loss plateaus.
A trained rLEM network can be used to autoregressively predict upstream phase space projections based on downstream phase space projections provided as inputs. The prediction procedure is detailed in Algorithm 1 and explained in Sec. 2.3. The algorithm operates with a user-defined downstream location (input) where the measurement is specified and the number of upstream locations where predictions are needed (output). It is flexible regarding inputs, accommodating multiple consecutive or non-consecutive downstream locations. Depending on the inputs, it can provide phase space projections for single, multiple, consecutive, or non-consecutive upstream locations.
Fig. 4 shows prediction in all the upstream modules given the last module (module 48) as the input. Similar plots using module 10 and module 25 as inputs are shown in Figs. A.3 and A.4. For brevity, three out of fifteen projections are shown across nine different modules. The original and predicted images can be compared using absolute difference plots. The mean squared error (MSE) and structural similarity index (SSIM) are also calculated and displayed alongside the absolute difference plots. It is evident from the figures and comparison metrics that rLEM performs well, demonstrating high accuracy prediction capabilities. We achieve an overall MSE and SSIM of ($\approx$ 5e-7, $\approx$ 0.998) and ($\approx$ 1e-6, $\approx$ 0.976) on the training and test set.
@@
Figure 4: Predictions from the trained rLEM at different upstream modules given module 48 as the input. The original, predicted and the absolute difference are plotted for (a) $x-y$ projection, (b) $E-\phi$ projection and (c) $x^{\prime}-y^{\prime}$ projection across nine different modules.
@@
Generating predictions for all 15 projections in all 47 modules takes less than one second with rLEM, whereas the simulator requires about 10 minutes on similar computing infrastructure. This results in a speedup factor of approximately 600.
3.3 Uncertainty analysis results
3.3.1 Aleatoric uncertainty in the latent space
@@
3.3.2 Aleatoric uncertainty propagation
@@ 4 Conclusions
CRediT authorship contribution statement
Mahindra Rautela: Conceptualization, Methodology, Data curation, Formal analysis, Investigation, Writing - original draft, Writing—review and editing. Alan Williams: Data curation, Investigation, Writing—review and editing. Alexander Scheinker: Conceptualization, Methodology, Formal analysis, Investigation, Writing—review and editing, Funding acquisition, Supervision.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
@@ Acknowledgements
This appendix contains supplementary information in the form of additional results from Sec. 3.
@@
Figure A.2: Comparison plot of two different random predictions ($A_{p}$/$A_{pred}$ and $B_{p}$/$B_{pred}$) from the trained rLEM at different upstream modules given two different module 48 as the input ($A_{t}$/$A_{true}$ and $B_{t}$/$B_{true}$): (a) $E-\phi$ projections (b) MSE vs module numbers.
@@
Figure A.3: Predictions from the trained rLEM at different upstream modules given module 10 as the input.
@@
Figure A.4: Predictions from the trained rLEM at different upstream modules given module 25 as the input.

### v2

@@ II.2 Reverse latent evolution model (RLEM)
Reversing the temporal dynamics to predict the current state $X_{t}$ given future states $X_{t+1},X_{t+2},\dots,X_{T}$ requires inverting the function $H$, which becomes a challenging inverse problem. Similar to [18], the inverse problem can be learned by calculating the joint probability distribution $P(X_{1},X_{2},...,X_{T})$ by factorising it using the chain rule of probability. The chain rule for probability is temporal direction-invariant, meaning it permits joint probability decomposition in either forward or reverse directions. Therefore, the joint probability distribution can be factored as in reverse direction as $P(X_{T})P(X_{T-1}|X_{T}),..,P(X_{0}|X_{1},X_{2},...,X_{T})$.
In particle accelerators, $X$ can represent phase space comprising of positions and momentum of billions of particles. However, the high-dimensionality of $X$ makes it difficult to calculate $P(X_{t}|X_{t+1},X_{t+2},\dots,X_{T})$. Fortunately, latent variable models (like variational autoencoders [63]) transform the higher-dimensional distribution $P(X_{1},X_{2},...,X_{T})$ into a lower-dimensional distribution $P(\mathbf{z}_{1},\mathbf{z}_{2},...,\mathbf{z}_{T})$, $p_{\theta}(\mathbf{z}|\mathbf{X})=p_{\theta}(\mathbf{X}|\mathbf{z})p(\mathbf{z})/p_{\theta}(\mathbf{X})$. The spatial dynamics is learned by minimizing the evidence lower-bound (ELBO) loss, $\mathbb{E}_{\mathbf{z}\sim q_{\phi}(\mathbf{z}|\mathbf{x},m)}[\log p_{\theta}(\mathbf{x}|\mathbf{z})]-D_{KL}(q_{\phi}(\mathbf{z}|\mathbf{x},m)||p_{\theta}(\mathbf{z}|m))$. The first term of the loss function captures the reconstruction error between the original and generated data, while the second term involves the Kullback–Leibler (KL) divergence, which quantifies the difference between the approximate and true posterior distributions [64].
In lower-dimensions, the temporal dynamics can be learned through the reverse autoregression $P(z_{t}|z_{t+1},z_{t+2},\dots,z_{T})$ to estimate $\mathbf{z}_{t}$. A recurrent neural network (RNN) like an LSTM can be used to invert the forward temporal dynamics in the latent space. The LSTM learns $P(z_{t}|z_{t+1},z_{t+2},\dots,z_{T})=f(z_{t}|h_{t+1},c_{t+1})$, where the hidden state, $h_{t+1}=g(z_{t+1},h_{t+2},c_{t+2})$ is calculated through future hidden and memory states and future observations [65]. A mean squared error based loss can be used to train the network $L(\psi;\mathbf{z})=\parallel\mathbf{z}_{t}-\hat{\mathbf{z}}_{t}\parallel_{2}$. Once the inverse temporal dynamics is learned in the latent space, the decoder part of the VAE transforms it back to the high-dimensional space, obtaining $P(X_{t}|X_{t+1},X_{t+2},\dots,X_{T})$.
@@
From a learning perspective, the autoregressive procedure used to train the LSTM is a crucial aspect of the latent evolution model. It is a more general formulation which does not assume a strict Markov property, i.e. $P(z_{t}|z_{t+1},z_{t+2},\dots,z_{T})=P(z_{t}|z_{t+1})$, on the temporal dynamics. Since the temporal dynamics are learned in a lower-dimensional latent space rather than the higher-dimensional image space ($n_{d}\approx 10^{6}$), $T$-step autoregression compensates for the information loss due to spatial dimensionality reduction. A single step autoregression model, using $P(z_{t}|z_{t+1})$, suffers in performance compared to the $T$-step autoregressive model. Additionally, autoregression procedure leverages temporal periodicity (short-term or long-term seasonality) present in the dataset better than Markov property.
II.2.1 RLEM for beam dynamics in LANSCE accelerator
The above mathematical explanations are more general and uses $t$ as a temporal index with $T$ as the final index where the states $X$ are measured or observed. We apply the RLEM to a dataset representing phase space (X) at 48 different accelerating modules of the LANSCE accelerator. Therefore, all the expressions can be written in terms of $m$, where $M=48$ is the final module. First, a conditional variational autoencoder (CVAE) projects the 15 unique projections of the 6D phase space of charged particle beams along with the module number into a lower-dimensional latent distribution. Subsequently, a LSTM network learns the reverse temporal dynamics within this latent space. The architecture of the RLEM is represented in Fig. 2.
@@ II.2.1 RLEM for beam dynamics in LANSCE accelerator
The model learns spatial and temporal correlations independently in two-step self-supervised process. The encoder performs convolution operations to extract features from the images. The module number is passed through dense layers and concatenated with the features vectors out of the encoder. The concatenated vector is the input to the latent space. The LSTM network operates on the learned latent space to predict the previous latent space points (corresponding to projections in the upstream modules) based on future points (corresponding to projections in the downstream modules). The predicted latent points are passed through a decoder which generates phase space projections across different modules of the accelerator. The methodology is outlined in Algorithm 1. The input of the algorithm is the user-defined phase space at downstream location of the measurement ($X_{m_{end}}$) and the upstream module/modules until where phase space predictions are desired, denoted by $m_{start}$. Typically, $m_{start}$ is set to 1, corresponding to the index of the first module, but it can also be any other module number $<m_{end}$. $m_{end}$ can be a single module, or multiple modules together, typically $m_{end}=M=48$. For the first iteration, the encoder of CVAE projects $X_{M:m}$ into the latent space to obtain $\mathbf{z}_{M:m}$. Subsequent iterations use only the LSTM and decoder of the CVAE. The LSTM acts on the latent representation to predict $\mathbf{z}_{m-1}$. The decoder of the CVAE then reconstructs $X_{m-1}$, providing the phase space projections at the previous module. The predicted latent point $\mathbf{z}_{m-1}$ is then concatenated with input $\mathbf{z}_{M:m}$ to form $\mathbf{z}_{M:m-1}$. This autoregressive procedure is iteratively applied to compute $\mathbf{z}_{m-2}$, followed by $X_{m-2}$, and so forth, until $m_{start}$.
Algorithm 1 Reverse latent evolution model (RLEM) to invert the beam dynamics
1:Inputs: State: {$X_{m}$,m}, Initial index: $m_{start}$, final index: $m_{end}$, pretrained CVAE model, pretrained LSTM model
2:for $m$ = $m_{end}$ to $m_{start}$ do
@@
4:         Compute latent encoding $\mathbf{z}_{M:m}=\text{CVAE-encoder}(X_{M:m},M:m)$
5:     Obtain $\mathbf{z}_{m-1}$ by passing $\mathbf{z}_{M:m}$ through the LSTM: $\mathbf{z}_{m-1}=\text{LSTM}(\mathbf{z}_{M:m})$
6:     Reconstruct $X_{m-1}$ by decoding $\mathbf{z}_{m-1}$: $X_{m-1}=\text{CVAE-decoder}(\mathbf{z}_{m-1})$
7:     Update latent sequence $\mathbf{z}_{M:m-1}$ by appending $\mathbf{z}_{m-1}$: $\mathbf{z}_{M:m-1}=\text{concat}(\mathbf{z}_{M:m},\mathbf{z}_{m-1})$
8:Return: reconstructed sequence $X_{m_{end}:m_{start}}$
III Results and Discussion
The training dataset comprising 1400 data objects is partitioned into training and validation sets with an 85:15 ratio. The encoder of the CVAE is composed of five convolutional layers with 32, 64, 128, 256, and 512 filters of size 3×3, each with a stride of 2, followed by a dense layer with 256 neurons. Additionally, a separate dense branch processes the module number through two layers of 32 neurons. The outputs of the primary and branch layers are concatenated and fed into the latent dense layer. We have performed a trial-and-error study balancing dimensions and loss to select the dimensionality of the latent space. This lower-dimensional representation enhances interpretability and reduces training time. The decoder mirrors the encoder’s architecture, featuring a dense layer followed by convolutional layers with 512, 256, 128, 64, and 32 filters of size 3×3. Each layer employs the LeakyReLU activation function and batch normalization.
@@ III Results and Discussion
III.1 Latent space visualizations
@@ III.2 Prediction results
The LSTM training with variable-length input and a single output allows the network to learn the autoregressive temporal patterns of the latent points, which is essential for making autoregressive predictions in the latent space. Our LSTM architecture comprises two layers, each with 64 hidden units. We minimize the mean squared error (MSE) between the true and predicted latent vectors using Adam optimizer. The initial learning rate is set to 0.001 and is halved after 10 epochs if the validation loss plateaus.
A trained RLEM network can be used to autoregressively predict upstream phase space projections based on downstream phase space projections provided as inputs. The prediction procedure is detailed in Algorithm LABEL:alg:predict and explained in Sec. II.2. The algorithm operates with a user-defined downstream location (input) where the measurement is specified and the number of upstream locations where predictions are needed (output). It is flexible regarding inputs, accommodating multiple consecutive or non-consecutive downstream locations. Depending on the inputs, it can provide phase space projections for single, multiple, consecutive, or non-consecutive upstream locations.
Fig. 4 shows prediction in all the upstream modules given the last module (module 48) as the input. Similar plots using module 10 and module 25 as inputs are shown in Figs. A.3 and A.4. For brevity, three out of fifteen projections are shown across nine different modules. The original and predicted images can be compared using absolute difference plots. The mean squared error (MSE) and structural similarity index (SSIM) are also calculated and displayed alongside the absolute difference plots. It is evident from the figures and comparison metrics that RLE performs well, demonstrating high accuracy prediction capabilities. We achieve an overall MSE and SSIM of ($\approx$ 5e-7, $\approx$ 0.998) and ($\approx$ 1e-6, $\approx$ 0.976) on the training and test set.
@@
Figure 4: Predictions from the trained RLE at different upstream modules given module 48 as the input. The original, predicted and the absolute difference are plotted for (a) $x-y$ projection, (b) $E-\phi$ projection and (c) $x^{\prime}-y^{\prime}$ projection across nine different modules.
@@
Generating predictions for all 15 projections in all 47 modules takes less than one second with RLE, whereas the simulator requires about 10 minutes on similar computing infrastructure. This results in a speedup factor of approximately 600.
III.3 Uncertainty analysis results
III.3.1 Aleatoric uncertainty in the latent space
@@
III.3.2 Aleatoric uncertainty propagation
@@ Acknowledgements
This appendix contains supplementary information in the form of additional results from Sec. III.
@@
Figure A.2: Comparison plot of two different random predictions ($A_{p}$/$A_{pred}$ and $B_{p}$/$B_{pred}$) from the trained RLE at different upstream modules given two different module 48 as the input ($A_{t}$/$A_{true}$ and $B_{t}$/$B_{true}$): (a) $E-\phi$ projections (b) MSE vs module numbers.
@@
Figure A.3: Predictions from the trained RLE at different upstream modules given module 10 as the input.
@@
Figure A.4: Predictions from the trained RLE at different upstream modules given module 25 as the input.